Ops/Comm Scheduling meeting : 19/12/2017¶
Location : CR21 at 14:00 pm
Phone access : phonebridge number
Access code: 1779339517
- MOM Update
- Data quality
- Detectors : Durga Run10581.pdf
- Target : Joe
- Physics shifter report: https://micewww.pp.rl.ac.uk/projects/analysis/wiki/2017-12-15_2017-12-19_physics-shifter
- RAL Christmas shutdown plans and supervision
- Prejudice is to shutdown as usual over Christmas and consider longer term plans in the New Year
- SS magnets and Focus coil
- Beamline conventional magnets
- Water cooling system
- Decay Solenoid
- Closed with mechanical frame installed.
- TOF/EMR/KL/CKOV ( gas / power )
- Trackers (gas / cryo / power )
- Electronics room
- January jobs:
- POMPOM tests 22nd-23rd Jan. (Henry)
- D2/D1 field mapping (Paolo)
- Earth Leakage measurements (Henry/Richard Hale)
- PPS Functional Test ???
- DSA PPS work (ISIS) 19th
- Synch closure 24th
- Chris Rogers' Decommissioning List
- SSU CC5 change?
- Channel 3T test?
- Next meeting - if required - on Jan 9
Updated by Nebrensky, Henry about 5 years ago · 22 revisions